A shoe store developed the following estimated regression equation relating sales to inventory investment and advertising expenditures.

ŷ = 29 + 12x1 + 7x2

where

x1 = inventory investment ($1,000s)
x2 = advertising expenditures ($1,000s)
y = sales ($1,000s).

(a) Predict the sales (in dollars) resulting from a $14,000 investment in inventory and an advertising budget of $11,000.

$

A shoe store developed the following estimated regression equation relating sales to inventory investment and advertising expenditures. ŷ = 27 + 12x1 + 6x2 where

x1 = inventory investment ($1,000s)
x2 = advertising expenditures ($1,000s)
y = sales ($1,000s).

Predict the sales (in dollars) resulting from a $14,000 investment in inventory and an advertising budget of $10,000.

$
Answer #1

The equation is given as:

ŷ = 29 + 12x1 + 7x2

So in order to predict sales for x1 = 14 and x2 = 11

So

ŷ = 29 + 12*14 + 7*11 = 274

So $ 274000 is the sales for the given inventory investment and advertising expenditures.

b)

Similarly for this equation :

ŷ = 27 + 12x1 + 6x2

x1 = 14

x2 = 10

ŷ = 27 + 12*14 + 6*10 = 255

So their will be $ 255000 sales for given inventory investment and advertising expenditures.